The p.d.f of a continuous random variable is give


P(x) ={ kx (1-x) e^x , 0 ≤ x ≤ 1

{ 0, otherwise

Find k and hence find mean and standard deviation


1
Expert's answer
2021-07-19T13:32:25-0400

"\int^1_0 kx(1-x)e^xdx=1"


"\int^1_0 kx(1-x)e^xdx=[ke^x(x-1)-ke^x(x^2-2x+2)]|^1_0="

"=-ke+k+2k"

"-ke+k+2k=1"

"k=\frac{1}{1-e}"


"E(X)=\int^1_0 kx^2(1-x)e^xdx=[ke^x(x^2-2x+2)-ke^x(x^3-3x^2+6x-6)]|^1_0="


"=3ke-2k-6k=\frac{3e-8}{1-e}"


"V(X)=E(X^2)-(E(X))^2"


"E(X^2)=\int^1_0 kx^3(1-x)e^xdx=[ke^x(x^3-3x^2+6x-6)-"


"-ke^x(x^4-4x^3+12x^2-24x+24)]|^1_0=-11ke+30k=\frac{30-11e}{1-e}"


"V(X)=\frac{30-11e}{1-e}-(\frac{3e-8}{1-e})^2=\frac{30-30e-11e+11e^2}{(1-e)^2}=\frac{11e^2-41e+30}{(1-e)^2}"


"\sigma=\sqrt{V(X)}=\frac{\sqrt{11e^2-41e+30}}{1-e}"